OpenStudy (anonymous):

What is the approximate volume of the cylinder? height = 52.2 cm; diameter = 20 cm A. 1640 cm3 B. 3300 cm3 C. 16,400 cm3 D. 65,600 cm3

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@babyboy2332

OpenStudy (johnweldon1993):

Well \[\large V = \pi r^2 h\] so \[\large V = \pi (10)^2 (52.2) = ?\]

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

well 100 times 52.2 = 522 so 522 times 3.14 would approximately be 1640